OpenStudy (mallorysipp234):

Simplify: (sin Θ − cos Θ)^2 − (sin Θ - cos Θ)^2

OpenStudy (solomonzelman):

a^2-a^2=0

OpenStudy (solomonzelman):

no trig needed.

OpenStudy (anonymous):

since here both brackets are same & we are subtracting them so it becomes 0 You can use this formula to expand brackets (a - b)^2 = a^2 - 2ab + b^2 , then simplify it

OpenStudy (anonymous):

so you will get 0